A restaurant owner is concerned about the amount of time customers must wait before being served over


weekends. He collects data on the waiting times, to the nearest minute, of 20 tables on Friday evening and


15 tables on Saturday evening.


The average waiting time on Friday evening is 24 minutes per table.


The total waiting time for the 15 tables on Saturday evening is seven hours and three


quarters of an hour.


The mean waiting time per table for Friday and Saturday evening, rounded to the nearest minute, is

1
Expert's answer
2022-04-03T14:42:25-0400

"7\frac{3}{4}h=465 min"

"465\/16=31" min for table on Saturday

15+20=35 tables on Friday and Saturday.

f(Friday)=20/35=0.57

f(Saturday)=15/35=0.43

"E(x)=\sum fx=0.57 \times24+ 0.43\times31=13.7+13.3=27"

Answer: 27 min